# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

### v2

@@ 1 Introduction
Time series forecasting is a cornerstone in diverse fields such as energy [23, 44], retail [9, 6], and finance [7], where accurate predictions can drive strategic decisions and enhance operational efficiency. In practical scenarios, forecasters have access not only to the target variables they aim to predict but also to covariates that offer rich contexts for predictions. For instance, the "Electricity Demand" data shown in Figure 1 (Left) demonstrates a strong negative correlation with the "Holiday" indicator, which is accessible up to the future forecasting horizon. This highlights the critical importance of appropriately incorporating covariates into the forecasting process.
@@
Figure 1: Left: Hourly electricity demand data for three weeks in Panama [2]. "Holiday" and "Weekend" indicators demonstrate a negative correlation with demand, providing important contextual information in the future forecasting horizon. "Humidity" displays a weak correlation but introduces temporal disturbances (red circle), which may distract the forecasting process. Right: Performance improvement of CITRAS through the utilization of past covariates and future covariates.
This practical problem setting presents two significant challenges. First, the forecasting model must be capable of universally accommodating heterogeneous variables. The availability and number of variables can vary depending on the scenario, and known covariates exhibit a length discrepancy compared to other variables. Second, the model must capture the dependencies between variables from both fine-grained and coarse-grained perspectives. The fine-grained perspective enables the model to precisely capture the local impact of covariates, such as sudden drops in electricity demand on holidays. However the model that relies solely on a fine-grained perspective may fail to capture global dependencies between variables, as it can overlook covariates representing infrequent events or can be distracted by temporal disturbances. Therefore, it is crucial for the model to capture dependencies between variables at both fine-grained and coarse-grained levels to ensure robust and accurate forecasting.
In recent years, deep forecasting models [42], particularly Transformers [40], have made remarkable progress. However, most Transformer-based models have primarily focused on multivariate settings that consider only multiple target variables. Some pioneering works [43, 30] have highlighted the importance of covariates; yet, the effective integration of heterogeneous covariates that leverages the inherent expressive power of Transformers requires further exploration. Existing approaches to capture cross-variate dependencies can be broadly categorized into two types: variate-level and patch-level. iTransformer [28] is a representative variate-level approach that captures inter-variate correlations by embedding each sequence into a single variate token and applying attention across multiple variate tokens. On the other hand, patch-level approaches, represented by Crossformer [55], applies cross-variate attention between locally semantic tokens obtained through patching. While variate-level approaches excel at capturing global dependency robustly against noisy interaction, patch-level approaches are better at capturing fine-grained information. Unfortunately, an architecture that enjoys the advantages of both of these approaches is yet to be developed.
@@ 2.1 Transformer-based Time Series Forecasting
Given the significant success of Transformers in the fields of natural language processing [12] and computer vision [15], numerous studies have attempted to leverage their capabilities for time series forecasting. These works can be roughly divided into two approaches: channel-dependent (CD) and channel-independent (CI). The CD approach assumes that the future values of a variable are determined by its past values as well as the values of other variables. In contrast, the CI approach posits that a specific variable only depends on its past values, omitting explicit interactions between variables.
@@
Early applications of Transformers in time series forecasting can be viewed as CD approaches, where multiple variables at the same time point are embedded into temporal tokens [56, 49, 57, 26, 13]. To address the computational costs and intricate dependencies arising from long time series data, Informer [56] designs a ProbSparse self-attention mechanism with a distilling operation to efficiently focus on important features. Autoformer [49] incorporates decomposition and an auto-correlation mechanism to uncover reliable dependencies from complex temporal patterns. However, point-wise temporal tokens have limited local semantics, making it difficult to capture intricate cross-time and cross-variate dependencies.
@@
In response to these challenges, PatchTST [34] adopts a CI approach and introduces patching, which handles time series data by segmenting it into fixed-length segments. Patching reduces computational costs and enhances the local semantics in token representation, making it a prevalent technique in subsequent methods [54, 53, 59]. Additionally, the CI approach is also adopted by many recent large models [32, 3, 11, 18] and LLM-based forecasters [19, 58, 20], as it avoids the risk of unnecessary noisy interactions between channels.
@@
However, the CI approach is sometimes regarded as an oversimplification [1], especially considering that fluctuations in time series data are often influenced by external factors. A representative CD approach is iTransformer [28], which embeds each channel into a single variate token and applies cross-variate attention over them. TimeXer [43] and Leddam [51] also capture cross-variate dependencies at the variate level, with the additional cross-time attention to capture intra-variate temporal dependencies. Meanwhile, Crossformer [55] applies cross-variate attention to patch-level representations. Any-variate Attention techniques [46, 30, 27] also apply attention to patch-level representations obtained by flattening all variables into one sequence. While the patch-level approach captures more fine-grained dependencies between variables, its local receptive field lacks global context awareness. Our method, CITRAS, first captures fine-grained dependencies at the patch level, and Attention Score Smoothing gradually refines them to achieve global variate-level dependencies.
@@ 2.2 Covariate-Informed Time Series Forecasting
To leverage the rich context offered by covariates, numerous approaches have been proposed, including statistical models [45, 33], MLP variants [10, 35, 8, 16], RNN variants [38], and Transformer-based models [24]. For instance, TiDE [10] introduces an MLP-based architecture that utilizes future covariates in the final projection to leverage their direct effects on future target variables. TSMixer-Ext [8] embeds past and future covariates into the same shape and applies feature mixing in the early stage. Temporal Fusion Transformer [24] designs gating mechanisms on top of the Transformer to select relevant features for forecasting. Contemporary to our work, ChronosX [4] introduces injection blocks to integrate covariates into pretrained models. A common drawback of these models is their structural complexity, which arises from architectural additions needed to handle heterogeneous variables.
Recent works, such as TimeXer [43] and Timer-XL [30], have attempted simpler integration of covariates into Transformers to retain their inherent expressive power. However, these approaches require further exploration. Specifically, the variate-level approach adopted by TimeXer is not compatible with decoder-only Transformers, leaving room for enhanced autoregressive capabilities through next-token prediction. The masking scheme in Timer-XL, known as TimeAttention, is not designed to accommodate future covariates, raising questions about its ability to utilize these pivotal features. In contrast to these works, CITRAS seamlessly integrates both observed and known covariates into the decoder-only Transformer, fully preserving its strong autoregressive capability.
@@ 3.2 Architecture
First, all sequences are segmented into non-overlapping patches of length $P$, and each patch is embedded into a $D$-dimensional token. For simplicity, we assume that both $T$ and $S$ are divisible by $P$. Taking target variables as an example, this can be formalized as:
$\{\mathbf{s}_{1}^{tgt,c},\mathbf{s}_{2}^{tgt,c},...,\mathbf{s}_{N_{tgt}}^{tgt,c}\}$ $=\operatorname{Patchfy}\left(\mathbf{X}_{1:T}^{tgt,c}\right),~{}~{}$ $c=1,...,C_{tgt}$ (2)
$\mathbf{H}_{i}^{tgt,c}$ $=\operatorname{Embed}\left(\mathbf{s}_{i}^{tgt,c}\right),~{}~{}$ $i=1,...,N_{tgt},~{}~{}c=1,...,C_{tgt}$
@@
In the cross-time attention module, we apply multi-head attention with causal masking to all variables to capture their intra-variate cross-time dependency. Here, we adopt Rotary Position Embedding (RoPE) [39] to capture temporal order. Taking target variables as an example, and dropping layer index for brevity, this can be formalized as:
$\widetilde{\mathbf{H}}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\mathbf{H}_{:}^{tgt,c}+\operatorname{MHA}\left(\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c}\right)\right),~{}~{}$ $c=1,...,C_{tgt}$ (3)
$\mathbf{H}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}\right)\right),~{}~{}$ $c=1,...,C_{tgt}$
where $\operatorname{LN}$ denotes layer normalization [5], $\operatorname{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ denotes the multi-head attention layer where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ serve as queries, keys and values, and $\operatorname{FFN}$ denotes a feed-forward network. Again, token embeddings of observed covariates $\mathbf{H}_{:}^{obs,c}$ and those of known covariates $\mathbf{H}_{:}^{knw,c}$ are similarly processed with the shared parameters.
@@
Attention is recognized as an effective approach for capturing cross-variate dependencies [28]. To represent the dependencies of target variables while avoiding unnecessary interactions among covariates, we adopt multi-head attention, where queries originate from the target variables, whereas keys and values are derived from all available variables. However, applying attention across the variables of the same patch steps, as common in [55, 1], fails to leverage pivotal future patches of known covariates, as they do not have corresponding target patches.
@@
[…]
$\widetilde{\mathbf{H}}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\mathbf{H}_{i}^{tgt,:}+\operatorname{MHA}\left(\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{k,:},\mathbf{H}_{i}^{v,:}\right)\right),~{}~{}$ $i=1,...,N_{tgt}$ (5)
$\mathbf{H}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}\right)\right),~{}~{}$ $i=1,...,N_{tgt}$
Intuitively, KV Shift allows the target variable (query) to incorporate the next-step future known covariate (value) based on the concurrent dependency on that covariate (key). As the target token gradually transforms into the prediction for the next step, this facilitates a natural flow of information for leveraging future known covariates.
@@
The attention score of $\mathrm{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ can be calculated as:
@@
To address this issue, we introduce Attention Score Smoothing in the multi-head attention layer of the cross-variate attention module, which transforms locally accurate patch-wise dependencies into global variate-level dependencies. Specifically, the attention score at each step is smoothed based on the attention scores up to that step in each head. We employ an Exponential Moving Average (EMA) as a smoothing method, as it enables adaptation to shifting correlations between variables [21] by assigning exponentially decreasing weights over time. Denoting original attention score at patch step $i$ as $\widetilde{\mathbf{A}}_{i}$, the smoothed attention score $\mathbf{A}_{i}$ can be calculated as:
$\mathbf{A}_{i}$ $=\alpha\widetilde{\mathbf{A}}_{i}+\left(1-\alpha\right)\mathbf{A}_{i-1},~{}~{}i=2,...,N_{tgt}$ (7)
where $\mathbf{A}_{1}=\widetilde{\mathbf{A}}_{1}$ and $\alpha$ is a smoothing factor, which we consider as a shared hyperparameter across all cross-variate attention modules and heads. It is important to note that the aim and structure of Attention Score Smoothing are completely different from existing integrations of EMA into attention mechanisms, as attempted in ETSformer [47] or CARD [41]. These works integrate EMA into cross-time attention to robustly capture intra-variate variations. In contrast, we apply EMA to cross-variate attention to refine cross-variate dependencies.
@@ 4.1 Covariate-Informed Forecasting
We use seven real world datasets for covariate-informed forecasting. (1)-(5) EPF [23]: Electricity Price Forecasting datasets of five major power markets spanning six years. The target variable is the hourly electricity price and we adopt day-ahead load forecasts and generation forecasts as known covariates. (6) EDF [2]: Electricity Demand Forecasting dataset in Panama spanning over five years. The target variable is the hourly national electricity demand in Panama. We adopt four meteorological indicators measured in three cities (total 12 indices) as observed covariates, and three binary calendar features (holiday, weekend, school day) as known covariates. (7) BS [17]: Bike Sharing dataset recording the hourly rental counts of Bikeshare system in Washington D.C. spanning two years. The target variables are three types of rental counts. We adopt five meteorological indicators as observed covariates and three calendar features (holiday, weekday, working day) as known covariates. Following previous works [35], we set the input length as 168 and forecasting length as 24. Further details about the datasets are included in Appendix A.1.
@@
We include the following baselines, categorized into four groups. (A) CD models that originally support future covariates, including TSMixer-Ext [8], TiDE [10], Temporal Fusion Transformer (TFT) [24], and TimeXer [43]. (B) CD models that are extended from original designs (as detailed in A.3) to support future covariates, including Timer-XL [30], iTransformer [28], and Leddam [51]. (C) CD models that cannot accommodate future covariates, including CARD [41], ModernTCN [14], TimesNet [48], and Crossformer [55]. (D) CI models including FITS [50], DLinear [52], and PatchTST [34]. For fair comparison with baselines (C) and (D), we also evaluate CITRAS without future covariates (i.e., without using KV Shift).
@@
Comprehensive results for covariate-informed forecasting are presented in Table 1. We observe that recent models capable of utilizing future covariates outperform those that cannot, underscoring the indispensability of future covariates in practical forecasting scenarios. Among these models, TimeXer captures cross-variate dependency at the variate level, while Timer-XL captures it at the patch level. The superiority of these contrasting methods is dataset-dependent. In contrast, CITRAS leverages the benefits of both approaches and seamlessly integrates future covariates, consistently demonstrating strong performance. Moreover, the superiority of CITRAS remains evident even when evaluated using conventional protocols that do not employ future covariates. This is because ordinary multivariate models treat target variables and covariates equally, leading to unnecessary interaction between covariates. In contrast, CITRAS applies attention only from target variables to covariates, avoiding this negative interaction.
Table 1: Full results of the covariate-informed forecasting with and without future covariates. The input length and forecasting length is set to 168 and 24, respectively. EPF results for "w/o future" are cited from TimeXer  [30] if available; other results are reproduced.
@@
4.2 Multivariate Forecasting
For a comprehensive comparison, we further evaluate our model on well-recognized multivariate forecasting datasets without using any covariates. To reflect practical scenarios, we adopt the rolling forecasting approach [29], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [29, 30], we set the input length to 672, the training forecasting length to 96, and the testing forecasting length to {96, 192, 336, 720}. We present the average results for the four forecasting lengths in Table 2. Compared to recent models specifically designed for these multivariate settings, CITRAS achieves superior or competitive performances in most datasets, highlighting the effectiveness and versatility of its design across various settings.
Table 2: Results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. We adopt the rolling forecasting approach  [29], where one model is used for four forecasting lengths. Averaged results are reported here and full results are provided in Table 10. Results are cited from Timer-XL [30] if available; otherwise reproduced.

### v3

@@ 1 Introduction
Time series forecasting is a cornerstone in diverse fields such as energy (Weron 2014), retail (Böse et al. 2017), and finance (Capistrán, Constandse, and Ramos-Francia 2010), where accurate predictions can drive strategic decisions and enhance operational efficiency. In practical scenarios, forecasters have access not only to the target variables they aim to predict but also to covariates that offer rich contexts for predictions. For instance, the ”Electricity Demand” data shown in Figure 1 (Left) demonstrates a strong negative correlation with the ”Holiday” indicator, which is accessible up to the future forecasting horizon. This highlights the critical importance of appropriately incorporating covariates into the forecasting process.
@@
Figure 1: Left: Hourly electricity demand data for three weeks in Panama (Aguilar Madrid and Antonio 2021). ”Holiday” and ”Weekend” indicators demonstrate a negative correlation with demand, providing important contextual information in the future forecasting horizon. ”Humidity” displays a weak correlation but introduces temporal disturbances (red circle), which may distract the forecasting process. Right: Performance improvement of CITRAS through the utilization of past covariates and future covariates.
This practical problem setting presents two significant challenges. First, the forecasting model must accommodate heterogeneous variables flexibly. The availability and number of variables can vary depending on the scenario, and known covariates exhibit a length discrepancy with other variables. Second, the model must capture the dependencies between variables from both fine-grained and coarse-grained perspectives. The fine-grained perspective enables the model to precisely capture the local impact of covariates, such as sudden drops in electricity demand on holidays. However, the model that relies solely on a fine-grained perspective may fail to capture global dependencies between variables, as it can overlook covariates representing infrequent events or can be distracted by temporal disturbances. Therefore, it is crucial for the model to capture dependencies between variables at both fine-grained and coarse-grained levels to ensure robust and accurate forecasting.
In recent years, deep forecasting models, particularly Transformers (Vaswani et al. 2017), have made remarkable progress. However, most Transformer-based models have primarily focused on multivariate settings that consider only multiple target variables. Some pioneering works (Wang et al. 2024b; Liu et al. 2025) have highlighted the importance of covariates; yet, the effective integration of heterogeneous covariates that leverages the inherent expressive power of Transformers requires further exploration. Existing approaches to capture cross-variate dependencies can be broadly categorized into two types: variate-level and patch-level. iTransformer (Liu et al. 2024b) is a representative variate-level approach that captures inter-variate correlations by embedding each sequence into a single variate token and applying attention across multiple variate tokens. On the other hand, patch-level approaches, represented by Crossformer (Zhang and Yan 2022), applies cross-variate attention between locally semantic tokens obtained through patching. While variate-level approaches excel at capturing global dependency robustly against noisy interaction, patch-level approaches are better at capturing fine-grained information. Unfortunately, an architecture that enjoys the advantages of both of these approaches is yet to be developed.
@@ 2.1 Transformer-based Time Series Forecasting
Given the significant success of Transformers in the fields of natural language processing (Devlin et al. 2018) and computer vision (Dosovitskiy et al. 2021), numerous studies have attempted to leverage their capabilities for time series forecasting. These works can be roughly divided into two approaches: channel-dependent (CD) and channel-independent (CI). The CD approach assumes that the future values of a variable are determined by its past values as well as the values of other variables. In contrast, the CI approach posits that a specific variable only depends on its past values, omitting explicit interactions between variables.
@@
Early applications of Transformers in time series forecasting can be viewed as CD approaches, where multiple variables at the same time point are embedded into temporal tokens (Zhou et al. 2022; Liu et al. 2022b; Dong et al. 2023). To address the computational costs and intricate dependencies arising from long time series data, Informer (Zhou et al. 2021) designs a ProbSparse self-attention mechanism with a distilling operation to efficiently focus on important features. Autoformer (Wu et al. 2021) incorporates decomposition and an auto-correlation mechanism to uncover reliable dependencies from complex temporal patterns. However, point-wise temporal tokens have limited local semantics, making it difficult to capture intricate cross-time and cross-variate dependencies.
@@
In response to these challenges, PatchTST (Nie et al. 2023) adopts a CI approach and introduces patching, which handles time series data by segmenting it into fixed-length segments. Patching reduces computational costs and enhances the local semantics in token representation, making it a prevalent technique in subsequent methods (Zhang et al. 2024b, a; Zhou et al. 2024). Additionally, the CI approach is also adopted by many recent large models (Ansari et al. 2024; Das et al. 2024; Goswami et al. 2024) and LLM-based forecasters (Gruver et al. 2023; Zhou et al. 2023; Jin et al. 2024), as it avoids the risk of unnecessary noisy interactions between channels.
@@
However, the CI approach is sometimes regarded as an oversimplification (Abdelmalak et al. 2025), especially considering that fluctuations in time series data are often influenced by external factors. A representative CD approach is iTransformer (Liu et al. 2024b), which embeds each channel into a single variate token and applies cross-variate attention over them. TimeXer (Wang et al. 2024b) and Leddam (Yu et al. 2024) also capture cross-variate dependencies at the variate level, with the additional cross-time attention to capture intra-variate temporal dependencies. Meanwhile, Crossformer (Zhang and Yan 2022) applies cross-variate attention to patch-level representations. Any-variate Attention techniques (Woo et al. 2024; Liu et al. 2025, 2024a) also apply attention to patch-level representations obtained by flattening all variables into one sequence. While the patch-level approach captures more fine-grained dependencies between variables, its local receptive field lacks global context awareness. Our method, CITRAS, first captures fine-grained dependencies at the patch level, and Attention Score Smoothing gradually refines them to achieve global variate-level dependencies.
@@ 2.2 Covariate-Informed Time Series Forecasting
To leverage the rich context offered by covariates, numerous approaches have been proposed, including statistical models (Williams 2001; Nicholson, Matteson, and Bien 2017), MLP variants (Das et al. 2023; Olivares et al. 2023; Chen et al. 2023; Ekambaram et al. 2024), RNN variants (Salinas et al. 2020), and Transformer-based models (Lim et al. 2021). For instance, TiDE (Das et al. 2023) introduces an MLP-based architecture that utilizes future covariates in the final projection to leverage their direct effects on future target variables. TSMixer-Ext (Chen et al. 2023) embeds past and future covariates into the same shape and applies feature mixing in the early stage. Temporal Fusion Transformer (Lim et al. 2021) designs gating mechanisms on top of the Transformer to select relevant features for forecasting. ChronosX (Arango et al. 2025) introduces injection blocks to integrate covariates into pretrained models. A common drawback of these models is their structural complexity, which arises from architectural additions needed to handle heterogeneous variables.
Recent works, such as TimeXer (Wang et al. 2024b) and Timer-XL (Liu et al. 2025), have attempted simpler integration of covariates into Transformers to retain their inherent expressive power. However, these approaches require further exploration. Specifically, the variate-level approach adopted by TimeXer is not compatible with decoder-only Transformers, leaving room for enhanced autoregressive capabilities through next-token prediction. The masking scheme in Timer-XL, known as TimeAttention, is not designed to accommodate future covariates, raising questions about its ability to utilize these pivotal features. In contrast to these works, CITRAS seamlessly integrates both observed and known covariates into the decoder-only Transformer, fully preserving its strong autoregressive capability.
@@ 3.2 Architecture
First, all sequences are segmented into non-overlapping patches of length $P$, and each patch is embedded into a $D$-dimensional token. For simplicity, we assume that both $T$ and $S$ are divisible by $P$. Taking target variables $c$ as an example, this can be formalized as:
$\{\mathbf{s}_{1}^{tgt,c},\mathbf{s}_{2}^{tgt,c},.,\mathbf{s}_{N_{tgt}}^{tgt,c}\}=\operatorname{Patchfy}\left(\mathbf{X}_{1:T}^{tgt,c}\right)$ (2)
$\mathbf{H}_{i}^{tgt,c}=\operatorname{Embed}\left(\mathbf{s}_{i}^{tgt,c}\right),~~i=1,.,N_{tgt}$
@@
In the cross-time attention module, we apply multi-head attention with causal masking to all variables to capture their intra-variate cross-time dependency. Following TimerXL (Liu et al. 2025), we adopt Rotary Position Embedding (RoPE) (Su et al. 2024) to capture temporal order. Taking target variables $c$ as an example, and dropping layer index for brevity, this can be formalized as:
$\widetilde{\mathbf{H}}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\mathbf{H}_{:}^{tgt,c}+\operatorname{MHA}\left(\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c}\right)\right)$ (3)
$\mathbf{H}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}\right)\right)$
where $\operatorname{LN}$ denotes layer normalization (Ba, Kiros, and Hinton 2016), $\operatorname{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ denotes the multi-head attention layer where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ serve as queries, keys and values, and $\operatorname{FFN}$ denotes a feed-forward network.
Again, token embeddings of observed covariates $\mathbf{H}_{:}^{obs,c}$ and those of known covariates $\mathbf{H}_{:}^{knw,c}$ are similarly processed with the shared parameters.
@@
Attention is recognized as an effective approach for capturing cross-variate dependencies (Liu et al. 2024b). To represent the dependencies of target variables while avoiding unnecessary interactions among covariates, we adopt multi-head attention, where queries originate from the target variables, whereas keys and values are derived from all available variables. However, applying attention across the variables of the same patch steps, as adopted by Crossformer (Zhang and Yan 2022), fails to leverage pivotal future patches of known covariates, as they do not have corresponding target patches.
@@
$\text{Key:}~~~\mathbf{H}_{i}^{k,:}$ $=\left[\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{obs,:},\mathbf{H}_{i}^{knw,:}\right]$ (4)
$\text{Value:}~~~\mathbf{H}_{i}^{v,:}$ $=\left[\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{obs,:},\mathbf{H}_{i+1}^{knw,:}\right]$
where $\left[\cdot,\cdot\right]$ denotes the concatenation along the variate dimension and $\mathbf{H}_{i}^{k,:},\mathbf{H}_{i}^{v,:}\in\mathbb{R}^{\left(C_{tgt}+C_{obs}+C_{knw}\right)\times D}$. After this, cross-variate dependency can be captured by the standard multi-head attention layer:
$\widetilde{\mathbf{H}}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\mathbf{H}_{i}^{tgt,:}+\operatorname{MHA}\left(\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{k,:},\mathbf{H}_{i}^{v,:}\right)\right)$ (5)
$\mathbf{H}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}\right)\right)$
for $i=1,...,N_{tgt}$. KV Shift maintains step alignment in the dot product calculation between queries and keys, enabling precise capture of concurrent dependencies. It then allows the target token to incorporate the next-step future known covariate (value) based on the strength of the obtained dependency. As the target token gradually transforms into the prediction for the next step, this facilitates a natural flow of information for leveraging future known covariates.
@@
The attention score of $\mathrm{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ is calculated as:
@@
To address this issue, we introduce Attention Score Smoothing in the multi-head attention layer of the cross-variate attention module, which transforms locally accurate patch-wise dependencies into global variate-level dependencies. Specifically, the attention score at each step is smoothed based on the attention scores up to that step in each head. We employ an Exponential Moving Average (EMA) as a smoothing method, as it enables adaptation to shifting correlations between variables (Kang, Shin, and Lee 2025) by assigning exponentially decreasing weights over time. Denoting original attention score at patch step $i$ as $\widetilde{\mathbf{A}}_{i}$, the smoothed attention score $\mathbf{A}_{i}$ can be calculated as:
$\mathbf{A}_{i}$ $=\alpha\widetilde{\mathbf{A}}_{i}+\left(1-\alpha\right)\mathbf{A}_{i-1},~~i=2,.,N_{tgt}$ (7)
where $\mathbf{A}_{1}=\widetilde{\mathbf{A}}_{1}$ and $\alpha$ is a smoothing factor, which we consider as a shared hyperparameter across all cross-variate attention modules and heads. It is important to note that the functionality provided by Attention Score Smoothing cannot be achieved by existing approaches that integrate EMA into attention mechanisms, as attempted in ETSformer (Woo et al. 2022) or CARD (Wang et al. 2024a). These works apply temporal smoothing in cross-time attention (i.e., along the parallel direction) to help capture temporal dependencies. In contrast, we apply temporal smoothing in cross-variate attention (i.e., along the orthogonal direction) to refine cross-variate dependencies over time.
@@ 4.1 Covariate-Informed Forecasting
We use seven real world datasets for covariate-informed forecasting. (1)-(5) EPF (Lago et al. 2021): Electricity Price Forecasting datasets of five major power markets spanning six years. The target variable is the hourly electricity price and we adopt day-ahead load forecasts and generation forecasts as known covariates. (6) EDF (Aguilar Madrid and Antonio 2021): Electricity Demand Forecasting dataset in Panama spanning over five years. The target variable is the hourly national electricity demand in Panama. We adopt four meteorological indicators measured in three cities (total 12 indices) as observed covariates, and three binary calendar features (holiday, weekend, school day) as known covariates. (7) BS (Fanaee-T 2013): Bike Sharing dataset recording the hourly rental counts of Bikeshare system in Washington D.C. spanning two years. The target variables are three types of rental counts. We adopt five meteorological indicators as observed covariates and three calendar features (holiday, weekday, working day) as known covariates. Following previous works (Olivares et al. 2023), we set the input length as 168 and forecasting length as 24. Further details about the datasets are included in Appendix A.1.
@@
We include the following baselines, categorized into four groups. (A) CD models that originally support future covariates, including TSMixer-Ext (Chen et al. 2023), TiDE (Das et al. 2023), Temporal Fusion Transformer (TFT) (Lim et al. 2021), and TimeXer (Wang et al. 2024b). (B) CD models that are extended from original designs (as detailed in A.3) to support future covariates, including Timer-XL (Liu et al. 2025), iTransformer (Liu et al. 2024b), and Leddam (Yu et al. 2024). (C) CD models that cannot accommodate future covariates, including CARD (Wang et al. 2024a), ModernTCN (Donghao and Xue 2024), TimesNet (Wu et al. 2023), and Crossformer (Zhang and Yan 2022). (D) CI models including FITS (Xu, Zeng, and Xu 2024), DLinear (Zeng et al. 2023), and PatchTST (Nie et al. 2023). For fair comparison with baselines (C) and (D), we also evaluate CITRAS without future covariates (i.e., without using KV Shift).
@@
Table 1 presents the results of covariate-informed forecasting with and without the use of future covariates. In the former setting, CITRAS outperforms TSMixer-Ext, TiDE, and TFT, all of which employ complex architectures to handle future covariates. This superior performance can be attributed to the simple yet effective design of CITRAS, which seamlessly integrates future covariates into the decoder-only Transformer through KV Shift, thereby preserving its inherent expressive power. Among recent models that adhere to the canonical Transformer architecture, TimeXer captures cross-variate dependencies at the variate level, while Timer-XL does so at the patch level. The relative superiority of these contrasting methods is dataset-dependent. In contrast, CITRAS leverages the advantages of both approaches through Attention Score Smoothing, consistently demonstrating strong performance across datasets. Moreover, the superiority of CITRAS remains evident even in conventional settings without future covariates. This is because standard multivariate models suffer from unnecessary interactions among covariates, as they treat target variables and covariates equally. In contrast, CITRAS applies attention only from target variables to covariates, thereby avoiding such negative interactions.
@@
Table 1: Full results of the covariate-informed forecasting with and without future covariates. The input length and forecasting length is set to 168 and 24, respectively. EPF results for ”w/o future” are cited from TimeXer  (Wang et al. 2024b) if available; other results are reproduced. The best result is represented in bold, followed by underline.
@@
Table 2: Results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. We adopt the rolling forecasting approach  (Liu et al. 2024c), where one model is used for four forecasting lengths. Averaged results are reported here and full results are provided in Table 8. Results are cited from Timer-XL (Liu et al. 2025) if available; otherwise reproduced.
4.2 Multivariate Forecasting
For a comprehensive comparison, we further evaluate our model on well-recognized multivariate forecasting datasets without using any covariates. To reflect practical scenarios, we adopt the rolling forecasting approach (Liu et al. 2024c), where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works (Liu et al. 2024c, 2025), we set the input length to 672, the training forecasting length to 96, and the testing forecasting length to {96, 192, 336, 720}. We present the average results for the four forecasting lengths in Table 2. Compared to recent models specifically designed for these multivariate settings, CITRAS achieves superior or competitive performances in most datasets, highlighting the effectiveness and versatility of its design across various settings.